Phytonadione, or more known as vitamin K1 , serves important roles in blood clotting mechanisms, bone health, and even bone health. When 0.368 g is dissolved in 12.0 g of camphor, the freezing point is lowered by 2.45 °C.



1
Expert's answer
2021-06-03T05:44:48-0400

Step 1: Calculate the freezing point depression of benzene.

Tf = (Freezing point of pure solvent) - (Freezing point of solution)

(5.5 oC) - (2.8 oC) = 2.7 oC

Step 2 : Calculate the molal concentration of the solution

molality = moles of solute / kg of solvent

moles of naphthalene = (1.60 g) (1 mol / 128 g) = 0.0125 mol naphthalene

molality of solution = (0.0125 mol) / (0.0200 kg) = 0.625 m

Step 3: Calculate Kf of the solution.

Tf = (Kf) (m)

(2.7 oC) = (Kf) (0.625 m)

Kf = 4.3 oC/m